OpenStudy (anonymous):

Find the solution ln(x-1)>1 for x>1

myininaya (myininaya):

x-1>e x>e+1

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

Nope thank you very much!!!!

OpenStudy (anonymous):

first change to experiential from \[e^{\ln(x-1)}>e ^{1}\] x-1>e x>e+1, since e+1 is always greater then 1, there is only one solution of x>e+1

OpenStudy (anonymous):

Thanks a lot,it is part of a function for me to solve and needen some help,!

myininaya (myininaya):

let me now if you need anymore help gl

OpenStudy (anonymous):

Just a note: if you have an inequality x > y, then\[e^x > e^y\]is also true because the exponential is a monotonically increasing function. This is not true for any function (for example,\[\sin(x) > \sin(y)\]may not be true).